For a school fund raiser, Jay sold 12 boxes of stationary adn 8 boxes of envelopes, for a total of $41.00. Sue sold 18 boxes of stationary and 16 boxes of envelopes, for a total of $68.50. How much did one box of envelopes cost?
Any help with this would be greatly appreciated!

Answer by Paul(988)

Let the satonary be x
Let the envelopes by y
12x+8y=41 --> solve for y
8y=41-12x
y=5.125-1.5x (subsititon)

18x+16y=68.50
Subsitute

18x+16(5.125-1.5x)=68.50
18x+82-24x=68.5
-6x=-13.5
x=2.25

y=5.125-1.5(2.25)
y=1.75

Hence, one box of envelop costs about $1.75.
Paul.